OpenStudy (anonymous):

Find the area inside the oval limaçon r = 6+sin θ

OpenStudy (anonymous):

as i recall it is \[\frac{1}{2}\int r^2 d\theta\]

OpenStudy (anonymous):

limits of integration from \(0\) to \(2\pi\)

OpenStudy (anonymous):

Hm, I'll try that. I seem to be using the wrong limits.

OpenStudy (anonymous):

since this is \(\sin(\theta)\) and not say \(\sin(2\theta)\) or some such thing, once complete rotation is from 0 to \(2\pi\)

OpenStudy (anonymous):

i think it is \(\frac{73\pi}{2}\) but i wouldn't bet money on it

OpenStudy (anonymous):

That's actually the correct answer - I've been sort of trying to backwards engineer it. I keep ending up with the wrong answer. I think I"m integrating wrong somehow. I keep ending up with \[37/2\theta - 6\cos \theta+(\sin2\theta)/4\]

OpenStudy (anonymous):

It looks like I'm dropping a 1/2 somewhere along the line - I think the formula should be \[73\theta/4 -6\cos \theta-1/8\sin2\theta \]

OpenStudy (anonymous):

\[\sin^2(\theta)=\frac{1}{2}(1-\cos(2\theta)\] so \[\int\sin^2(\theta)=\frac{\theta}{2}-\frac{1}{4}\sin(2\theta)\]

OpenStudy (anonymous):

\[36+12\sin(\theta)+\sin^2(\theta)\] anti derivative is \[36\theta-12\cos(\theta)+\frac{\theta}{2}-\frac{1}{4}\sin(2\theta)\]

OpenStudy (anonymous):

\[1/2\int\limits_{0}^{2\pi}(6+\sin)^2\] \[=1/2\int\limits_{0}^{2\pi}36+12\sin \theta+\sin^2\] \[= 1/2(36\theta - 12\cos \theta +\int\limits \sin^2\theta)\] \[=1/2 (36\theta - 12\cos \theta+\int\limits (1-\cos2\theta)/2\] \[= 1/2(36\theta-12\cos \theta+\theta/2-(\sin2\theta))/2\] \[=37/2\theta-6\cos \theta-(\sin2\theta)/4\]

OpenStudy (anonymous):

when you plug in 0 you get -12 when you evaluate at \(2\pi\) you get \[72\pi+\pi-12\]

OpenStudy (anonymous):

subtract and get \(73\pi\) then divide by 2

OpenStudy (anonymous):

oh my god, I have been forgetting to multiply the 36 by 2 before adding it to the 1/2.

OpenStudy (anonymous):

Thank you so much for all your help - I've been working on this one problem for hours.